# Legal: Enhancing Legal Reasoning in LLMs via Reinforcement Learning with Chain-of-Thought Guided Information Gain

### v1

@@ 5.2 Ablation Study
To systematically assess the contribution of each core component in the Legal framework, we conduct a series of ablation experiments. The results are summarized in Table 2.
@@
Among all components, the information-gain-enhanced reward mechanism contributes most significantly to performance improvements. To verify this, we evaluate a variant that excludes information gain and instead relies solely on traditional reward signals (w/o Info-Gain). The resulting performance degradation confirms the value of incorporating differential Q-value analysis, which enables fine-grained differentiation of reasoning quality. The gap between this variant and the full Legal model underscores the effectiveness of our reward enhancement in capturing the subtle characteristics of high-quality legal reasoning.
@@
| Qwen2.5-7B-Instruct | ||||||
| Zero-shot | 77.10 | 53.81 | 48.30 | 59.80 | 66.80 | 61.16 |
| Legal | 83.14 | 68.53 | 64.80 | 64.80 | 75.80 | 71.41 |
| w/o SFT | 84.11 | 62.82 | 72.10 | 62.20 | 69.80 | 70.21 |
| w/o GRPO | 80.14 | 54.73 | 50.20 | 55.60 | 69.20 | 61.97 |
| w/o Info-Gain | 84.82 | 62.07 | 61.40 | 60.60 | 72.00 | 68.18 |
| Qwen2.5-14B-Instruct | ||||||
| Zero-shot | 84.18 | 58.50 | 42.60 | 55.40 | 76.80 | 63.50 |
@@
5.3 The Effectiveness of Information-Gain Enhanced Reward Modeling
@@
In this section, we analyze the effectiveness of the information gain module during training through two perspectives: (1) its influence on model confidence via logits analysis, and (2) its impact on training efficiency, measured by batch reward variance.
@@
In deep learning, logits often reflect the model’s confidence in its predictions. Higher logits indicate stronger confidence, resulting from more stable and consistent internal representations, which manifest as concentrated probability distributions. To evaluate confidence improvement, we compare the logits performance of two variants of Legal, depending on whether the information gain module is incorporated. As demonstrated in Figure LABEL:fig:Logits_compare, in direct response scenarios (Figure LABEL:fig:logits_a), both models exhibit similar performance during early training stages. However, Legal with information gain module progressively develops higher confidence levels, with increasingly pronounced performance gaps in later stages. This improvement becomes dramatically more evident in Chain-of-Thought prompting scenarios (Figure LABEL:fig:logits_b), where Legal demonstrates substantial and consistent logits gains throughout training while the baseline variant maintains relatively low confidence levels.
The improved logit values confirm that our information gain mechanism successfully identifies high-quality reasoning patterns. This is further supported by the positive correlation between information gain and legal reward, as illustrated in Figure LABEL:fig:reward_a, where both information gain and reward values demonstrate consistent growth across training stages. This correlation indicates that samples with higher information gain consistently receive higher rewards, showing that information gain provides another way to evaluate the quality of generated responses during GRPO training. Based on this positive correlation, our mechanism produces significant improvements in training efficiency and stability. As shown in Figure LABEL:fig:reward_b, Legal with information gain enhancement maintains more stable training dynamics with substantially reduced reward variance compared to the baseline variant, creating more consistent reward signals aligned with reasoning quality.
5.4 Performance Quality Assessment
This experiment further explores the effectiveness of Legal through performance quality assessment from both token-level and sequence-level perspectives.
Token-Level Analyses
We first analyze the prominence of legal-related tokens. Legal token prominence, defined as the difference between logit values for specific legal tokens and the average logit value across the entire sequence, serves as an indicator of the model’s focus on legally relevant content.
@@ 5.4 Performance Quality Assessment
As shown in Figure LABEL:fig:token_a, we observe a progressive improvement across training stages: starting from the Zero-shot baseline with mainly negative prominence values, R1-Distill demonstrates initial improvements through high-quality reasoning data distilled from a large reasoning model, establishing foundational capabilities for identifying legally relevant content. Legal further advances this progression through GRPO training, with the prominence values becoming notably higher and more samples achieving positive values compared to previous stages, indicating substantially enhanced attention to legally relevant tokens.
@@
We further assess the quality of reasoning chains generated by our models through perplexity analysis for ground truth prediction. Specifically, we compare Legal with and without information gain enhancement by asking each model to generate reasoning chains for questions and then feeding these reasoning chains along with questions to the vanilla LLM (Qwen2.5-32B-Instruct) to calculate the perplexity of ground truth answers. Lower perplexity indicates that the reasoning chain better supports the correct answer prediction, reflecting higher reasoning quality. As demonstrated in Figure LABEL:fig:token_b, Legal with information gain enhancement consistently achieves lower perplexity values across legal reasoning tasks, with statistically significant improvements confirmed by paired t-test results. This demonstrates that the information gain module effectively enhances the quality of generated reasoning chains, enabling the LLM to generate more accurate responses.
| Model | Short Reasoning | Long Reasoning | ||
| Length | Average Score | Length | Average Score | |
| Legal | 222 | 77.41 | 299 | 73.69 |
| Legal w/o Info-Gain | 251 | 76.36 | 341 | 69.10 |
Sequence-Level Analyses
We further examine how the addition of information gain module affects reasoning quality across different scenario complexities. We categorize evaluation samples based on the inherent complexity of legal reasoning tasks, dividing them into short and long scenarios according to the response length generated by the baseline model Legal w/o Info-Gain. As shown in Table 3, Legal, which incorporates the information gain module consistently outperforms the baseline model across both short and long scenarios, generating more concise responses with superior performance scores. Notably, the performance improvement margin from adding the information gain module is more pronounced in long scenario, demonstrating that this incremental enhancement is particularly effective for legal reasoning tasks that require longer chain-of-thought.
@@ 6 Conclusions
This paper proposes Legal, a novel framework that encourages the model to explore diverse reasoning trajectories and develop robust reasoning strategies by incorporating chain-of-thought-guided information gain into Reinforcement Learning with Verifiable Rewards (RLVR). Experimental results demonstrate that Legal effectively balances direct-answer and reasoning-enhanced generation modes, enabling the model to produce more concise and accurate responses. Further analysis shows that the reasoning processes generated by our method exhibit superior quality compared to baseline models, yielding significant performance gains.
Appendix A Appendix
A.1 Dataset Statistics
As shown in Table 4, we use datasets from CAIL2018, JEC_QA, and LAIC2021 for training, covering five legal reasoning tasks. For evaluation, we conduct both in-domain evaluation using LawBench and LexEval benchmarks, and out-of-domain evaluation using six sub-tasks from the DiscLaw benchmark to assess model generalization capability.
A.2 Training Details
For Legal, we apply LoRA to the query, key, value, output, gate, up, and down projection matrices. The maximum prompt length is set to tokens while maximum completion length is set to tokens. We train for epoch for the 14B model and epochs for the other size models, using the cosine learning rate scheduler with a warmup ratio of . For optimization, we employ the 8-bit AdamW optimizer with , , and weight decay of . We set the maximum gradient norm to .
A.3 Legal Reward Construction
We design three distinct reward mechanisms based on the nature of expected outputs and evaluation requirements:
F1-Score Based Reward: Provides balanced evaluation for multi-label scenarios:
This mechanism effectively penalizes both incomplete responses and excessive false positive predictions.
Distance-Based Reward: Offers continuous evaluation for numerical predictions:
where is a normalization constant set to 36 in our experiments. This design provides partial credit for predictions approximating the correct range.
Exact Match Reward: Provides strict binary evaluation for categorical selections:
These mechanisms are strategically applied: for criminal charge prediction and legal article identification, for financial amount calculation and multiple choice questions, and for sentencing prediction with wide value ranges.
| Task | Dataset | Metric | Quantity |
| Training | |||
| Article Prediction | CAIL2018 | F1 | 1,396 / 150 |
| Cause Prediction | CAIL2018 | F1 | 1,373 / 169 |
| Penalty Prediction | CAIL2018 | Distance | 1,402 / 144 |
| Case Analysis | JEC_QA | Accuracy | 1,403 / 50 |
| Financial Calculation | LAIC2021 | Accuracy | 1,407 / 50 |
| In-domain Evaluation | |||
| Article Prediction | CAIL2018 | F1 | 500 |
| Cause Prediction | CAIL2018 | F1 | 500 |
| Penalty Prediction | CAIL2018 | Accuracy | 1,000 |
| Case Analysis | JEC_QA | Accuracy | 500 |
| Financial Calculation | LAIC2021 | Accuracy | 500 |
| Out-of-domain Evaluation | |||
| National Judicial Exam | NJE | Accuracy | 537 |
| Patent Agent Exam | PAE | Accuracy | 118 |
| Graduate Entrance Exam | UNGEE | Accuracy | 320 |
| Legal Basic Knowledge | LBK | Accuracy | 275 |
| Public Institutions Exam | PFE | Accuracy | 170 |
A.4 Proof of Logits-Q Function Equivalence
Given dataset , the next-token prediction objective is:
while the offline inverse reinforcement learning objective is:
Since , Eq. (14) can be rewritten as:
| Case Analysis Task on Cross-border Civil Procedure | |
| Criminal Fact | Dayang Company, with its domicile in District A, City A of China, signed a sales contract with Kerrys Company in Germany. The German Kerrys Company has an office established in China, and later the parties had a dispute over quality issues. Regarding statements about transnational civil litigation, which of the following options is INCORRECT? A: Both parties may agree to determine the court with jurisdiction over the dispute B: The court may serve litigation documents to Kerrys Company’s office in China C: Both parties have an appeal period of 30 days D: Kerrys Company may appoint a lawyer from its home country to serve as a litigation representative in a non-lawyer capacity |
| Vanilla | reasoningOption A Analysis: According to Article 34 of the Civil Procedure Law of the People’s Republic of China, parties to contract or other property rights disputes may agree in writing to choose the people’s court with jurisdiction at the defendant’s domicile, contract performance location, contract signing location, plaintiff’s domicile, or location of the subject matter that has actual connection with the dispute. Therefore, both parties can agree to determine the court with jurisdiction over the dispute. Option B Analysis: According to Article 265 of the Civil Procedure Law of the People’s Republic of China, in transnational civil litigation, if the law of the country where the recipient is located allows service by mail, service by mail is permitted. If no service receipt is returned within three months from the date of mailing, but various circumstances sufficiently indicate delivery has occurred, it shall be deemed served upon expiration of the period. Since Kerrys Company has an office in China, the court can serve litigation documents to Kerrys Company’s office. Option C Analysis: According to Article 267 of the Civil Procedure Law of the People’s Republic of China, in transnational civil litigation, the appeal period is 30 days. Therefore, both parties have an appeal period of 30 days. Option D Analysis: According to Article 298 of the Civil Procedure Law of the People’s Republic of China, when foreigners, stateless persons, foreign enterprises and organizations file lawsuits or respond to lawsuits in people’s courts and need to appoint lawyers for litigation representation, they must appoint lawyers from the People’s Republic of China. Therefore, Kerrys Company cannot appoint a lawyer from its home country to serve as a litigation representative in a non-lawyer capacity. Reasons for Excluding Options: Options A, B, and C all comply with legal provisions, while Option D does not comply with legal provisions. Correct Option Confirmation: Option D does not comply with legal provisions, therefore it is the incorrect option. /reasoning answer D /answer |
| Legal | reasoningOption A Analysis: Correct, both parties may agree to determine the court with jurisdiction. Option B Analysis: Correct, the court may serve documents to foreign parties at their offices in China. Option C Analysis: Incorrect, Chinese parties have an appeal period of 15 days, foreign parties have 30 days. Option D Analysis: Correct, foreign parties may appoint lawyers from their home country to serve in a non-lawyer capacity as representatives. Reasons for Excluding Options: Option C - Chinese citizens have 15 days, foreign citizens have 30 days. Correct Option Confirmation: C /reasoning answer C /answer |
A.5 Case Study
In Table 5, we compare the performance of Legal and zero-shot models in analyzing a cross-border civil procedure case involving a commercial dispute between Chinese and German companies. The zero-shot model exhibits several critical limitations that significantly impair its practical utility. The model not only conflates fundamental procedural distinctions between domestic and international litigants—particularly by erroneously equating the 15-day appeal period for Chinese parties with the 30-day requirement for foreign-related cases—but also fails to recognize valid exceptions in litigation representation rules. These shortcomings are further exacerbated by the model’s generation of verbose and unfocused legal citations, which obscures crucial jurisdictional differences.
Legal demonstrates markedly superior performance through its nuanced understanding of cross-border procedural law. The model accurately identifies and explains key asymmetries in legal treatment, including differentiated appeal periods and permissible representation forms. Its analytical approach represents a significant technical advancement, employing structured reasoning to strategically highlight legally determinative factors while maintaining clear and concise arguments. This sophisticated methodology stems from specialized training on party-specific rules and exceptions in international litigation. The model’s question-oriented output mechanism ensures direct relevance to the core legal issues, avoiding the digressive tendencies characteristic of zero-shot model responses. This demonstrates that Legal possesses stronger legal reasoning capabilities, enabling it to derive correct judgments and conclusions through key facts.
A.6 Prompt Template
In this section, we introduce the prompt used in Legal. Figure 4 presents the prompt used in direct mode. Figure 5 indicates the prompt used with Chain of Thought during the rollout process.

### v3

@@ 4.2 Ablation Study
To systematically assess the contribution of each core component in the Legal framework, we conduct ablation experiment on Qwen2.5-14B-Instruct. The results are summarized in Table 2.
@@
The information-gain-enhanced reward mechanism also provides notable performance gains. To verify this, we evaluate a variant that excludes information gain and instead relies solely on traditional reward signals (w/o Info-Gain). The resulting performance degradation confirms the important role of incorporating differential Q-value analysis, which enables fine-grained differentiation of reasoning quality. The performance gap between the baseline variant that excludes information gain and our full Legal model demonstrates that our information-gain enhancement effectively improves legal reasoning ability.
@@
| Zero-shot | 84.18 | 58.50 | 42.60 | 55.40 | 76.80 | 63.50 |
4.3 Performance of Information-Gain Enhanced Reward
@@
In this section, we analyze the effectiveness of the information gain reward modeling from two perspectives: (1) its influence on model confidence via logits analysis during training, and (2) its effectiveness on token-level performance quality assessment.
@@
Logits Analysis and Model Confidence. As demonstrated in Figure LABEL:fig:logits_c and Figure LABEL:fig:logits_d, the figures illustrate how average logits change with training steps for the direct response and CoT prompting scenarios, respectively. To evaluate how the information gain reward modeling contributes to model confidence, we compare the logits performance of two Legal variants, depending on whether information gain reward is included. Logits reflect model confidence in predictions and higher values indicate stronger confidence from more stable internal representations. In direct response scenario, both variants perform similarly initially, but our full Legal progressively develops higher confidence with widening performance gaps in later training stages. This improvement is more pronounced in CoT prompting scenario, where our Legal shows substantial and consistent logits gains throughout training while the other variant maintains relatively low confidence levels. These improved logits confirm that our information gain mechanism effectively identifies high-quality reasoning patterns.
@@
Performance Quality Assessment. As shown in Figure LABEL:fig:token_a, we evaluate legal token prominence across three different models. Legal token prominence is defined as the difference between logit values for specific legal tokens and the average logits across the sequence, which measures the model’s attention to legally relevant content. We observe progressive improvements: from negative baseline values in Zero-shot, R1-Distill achieves initial improvements through distilled reasoning data, while Legal further advances through GRPO training with notably higher prominence values and more positive samples, indicating enhanced attention to legal tokens.
@@
We further assess reasoning quality through perplexity analysis by comparing Legal with and without information gain enhancement in Figure LABEL:fig:token_b. For each question, we generate reasoning chains and provide them, along with the question, to a vanilla LLM (Qwen2.5-32B-Instruct) to calculate ground truth answer perplexity. Legal with information gain consistently achieves lower perplexity values across tasks, demonstrating the information-gain reward modeling helps to conduct higher quality reasoning results.
@@ 5 Conclusions
This paper proposes Legal, a novel framework that encourages LLMs to explore diverse reasoning trajectories and develop robust reasoning strategies for legal tasks by incorporating chain-of-thought–guided information gain into Reinforcement Learning with Verifiable Rewards (RLVR). Experimental results demonstrate that Legal effectively balances direct-answer and reasoning-enhanced generation modes, improving model confidence, emphasizing legally relevant tokens and enhancing the quality of the reasoning process compared to baselines.
Acknowledgments
This work is partly supported by the National Natural Science Foundation of China (No. 62576082 and No. 62461146205).